# Asymptotic directions in the moduli space of curves

### v1

@@ 1. Introduction
Clearly a tangent direction to a totally geodesic subvariety is asymptotic. But the locus of asymptotic directions in the projective tangent bundle of the moduli space of curves ${\mathcal{M}}_{g}$ is a natural locus, that is worthwhile investigating. For example we believe that asymptotic directions could also be useful in the study of fibred surfaces in relation with Xiao conjecture ([29], [13]).
Since $II$ is injective (see [4, Corollary 3.4]), its image in $Sym^{2}H^{1}(T_{C})^{\vee}$ is a linear system of quadrics in ${\mathbb{P}}H^{1}(T_{C})\cong{\mathbb{P}}^{3g-4}$ of dimension $\frac{(g-2)(g-3)}{2}$. Hence for every curve $C$ of genus $g\leq 9$, $dim(II(I_{2}))<3g-4$, so the intersections of the quadrics in $II(I_{2})$ is non empty, thus there exist asymptotic directions. In fact for $g\leq 7$ there are examples of special subvarieties of ${\mathcal{A}}_{g}$ generically contained in the Torelli locus for $g\leq 7$ (see [7], [8], [9], [10], [22], [23], [24], [26], [27]).
@@
(Theorem 5.5, Theorem 5.1) Let $C$ be a smooth curve of genus $g$, take an integer $d<Cliff(C)$ and an infinitesimal deformation $\zeta\in H^{1}(T_{C})$ of rank $d$. Then we have:
@@
as the locus of deformations of rank $d$. When $d=1$ it is the bicanonical curve.
Since the Clifford index of the general curve of genus $g$ is $\lfloor\frac{g-1}{2}\rfloor$, this gives a characterisation of the image of the natural map ${\mathbb{P}}T_{C_{d}}\to{\mathbb{P}}(H^{1}(T_{C}))$ as the locus of deformations of rank $d$, for all $d<\lfloor\frac{g-1}{2}\rfloor$. Moreover, for such values of $d$, Theorem 1.1(2) says that for the general curve of genus $g$ the base locus of the linear system of quadrics $II(I_{2})$ in ${\mathbb{P}}H^{1}(T_{C})$, does not contain any point $[\zeta]$ with with $Rank(\zeta)=d$.
@@
More precisely, the linear combinations of Schiffer variations which are asymptotic in the above Theorem are $\xi_{p}\pm i\xi_{\sigma(p)}$, where $\sigma$ is the bielliptic involution and $(p,\sigma(p))\in C\times C$ is in the zero locus of the meromorphic form $\hat{\eta}\in H^{0}(K_{C\times C}(2\Delta))$ which determines the second fundamental form $II$ (see [3, Theorem 3.7]).
@@
In Section 4 we recall the definition of the second fundamental form, we define asymptotic directions and we use the results in Section 3 to give a formula that computes the second fundamental form $II(Q)(\zeta\odot\zeta)$ on some $\zeta$ with non trivial kernel, in terms or residues of meromorphic forms (see Proposition 4.3).
In Section 5 we consider infinitesimal deformations of rank $d$ less than the Clifford index of the curve, and we analyse the extension corresponding to $\zeta$ and the rank 2 vector bundle $E$ of the extension. First we show that if either $d<Cliff(C)$, or $d=Cliff(C)<\frac{g-1}{2}$ and $E$ is not globally generated, then $\zeta$ is a linear combination of Schiffer variations supported on a divisor $D$ of degree $d$ (see Theorem 5.1). One of the main technical tool is a Theorem of Segre-Nagata and Ghione (see [19] p. 84) on the existence of a subline bundle $A$ of $E$ such that $deg(A)\geq\frac{g-1}{2}$. Then we show that if $d<Cliff(C)$, no linear combinations of Schiffer variations of rank $d$ is asymptotic (see Theorem 5.4). Here we use a result of Green and Lazarsfeld ([15, Theorem 1]) that allows us to find some quadrics where our technique to compute the second fundamental form works well. Finally we prove Theorem 5.5.
@@ 2. Preliminaries on deformations
Let $s\in H^{0}({\mathcal{O}}_{Z})$ be a holomorphic section and assume that $f_{i}(z_{i})=\sum_{j=0}^{n_{i}}\beta_{j,i}z_{i}^{j}$ is its polynomial expression in $U_{i}$. Let $\tilde{\rho}$ be a ${\mathcal{C}}^{\infty}$ function on $C$ which is equal to $1$ on $D_{i}$ and equal to 00 on $C\setminus\cup_{i}\Delta_{i}.$
@@
Then we set in Dolbeault cohomology $\zeta:=[{\bar{\partial}}(\frac{\rho}{\omega})]=\delta(s).$ One clearly has $\omega\cup\zeta=0$ and conversely any $\zeta$ such that $\omega\cup\zeta=0$ has a Dolbeault representative as above.
@@
We say that an infinitesimal deformation $\zeta\in H^{1}(T_{C})$ is split if the vector bundle $E$ in (1) splits as a direct sum of line bundles $E=(K_{C}\otimes L^{\vee})\oplus L$.
@@
Let $C$ be a smooth complex projective curve of genus $g\geq 2$. Let $p$ be a point in $C$ and $z$ a local coordinate centred in $p$. For $31\leq n\leq 3g-$ 3, we define the $n^{th}$ Schiffer variation at $p$ to be the element $\xi_{p}^{n}\in H^{1}(C,T_{C})\cong H^{0,1}_{\bar{\partial}}(T_{C})$ whose Dolbeault representative is ¯$\frac{\bar{\partial}\rho_{p}}{z^{n}}\frac{\partial}{\partial z}$, where $\rho_{p}$ is a bump function in $p$ which is equal to one in a small neighborhood $U$ containing $p$, $\xi_{p}^{n}=[\frac{\bar{\partial}\rho_{p}}{z^{n}}\frac{\partial}{\partial z}]$. Clearly $\xi_{p}^{n}$ depends on the choice of the local coordinate $z$. Take $31\leq n\leq 3g-$ 3. Take the exact sequence
@@
In particular, any linear combination $\zeta=\sum_{i=1}^{k}\sum_{j=1}^{m_{i}}b_{j,i}\xi^{j}_{p_{i}}$, with $\sum_{i=1}^{k}m_{i}\leq g$ has rank $\leq\sum_{i=1}^{k}m_{i}$, since $H^{0}(K_{C}(-\sum_{i=1}^{k}m_{i}p_{i}))\subset ker(\cup\zeta)$.
@@ 3. Some computations
Consider two holomorphic one forms $\omega_{1}$ and $\omega_{2}$. With the notation introduced in section 2, take an infinitesimal deformation $\zeta$ having a form $\omega\neq 0$ in its kernel and denote by $Z=\sum n_{i}p_{i}$ its zero divisor . Then $\zeta=[\Theta]\in H^{1}(T_{C})$ with $\Theta:={\bar{\partial}}(\frac{\rho}{\omega})$, and $\rho=\sum_{i}\tilde{\rho}f_{i}$. We have $\Theta\omega_{i}=\gamma_{i}+{\bar{\partial}}h_{i}$, where $h_{i}$ is a ${\mathcal{C}}^{\infty}$ function on $C$, so $\zeta\cup\omega_{i}=[\gamma_{i}]\in H^{1}({\mathcal{O}}_{C})$. Consider the two meromorphic functions $g_{1}=\frac{\omega_{1}}{\omega}$ and $g_{2}=\frac{\omega_{2}}{\omega}$. Then clearly, by the above construction we have
@@
Note that if $\rho g_{1}$ is $C^{\infty}$, up to a constant we have $h_{1}=\rho g_{1},$ $\gamma_{1}=0$, hence $\omega_{1}\in ker(\cup\zeta)$.
@@
Since $\int_{C}\partial h_{1}\wedge\gamma_{2}=\int_{C}d(h_{1}\gamma_{2})=0$, we get
@@
$=\int_{C}\partial(\rho g_{1})\wedge{\bar{\partial}}(\rho g_{2})=\int_{C}d(\rho g_{1})\wedge{\bar{\partial}}(\rho g_{2})=\int_{C}d(\rho g_{1}{\bar{\partial}}(\rho g_{2}))-\int_{C}\rho g_{1}\partial({\bar{\partial}}(\rho g_{2})).$
Now,
$\int_{C}d(\rho g_{1}{\bar{\partial}}(\rho g_{2}))=\sum\lim_{\epsilon\to 0}\int_{\partial D_{i,\epsilon}}\rho g_{1}g_{2}{\bar{\partial}}\rho=0$
since ¯${\bar{\partial}}\rho=0$ near the points $p_{i}.$ Here $D_{i,\epsilon}$ is a small disc around $p_{i}$. It follows then
$\int_{C}\partial h_{1}\wedge{\bar{\partial}}h_{2}=-\int_{C}\rho g_{1}\partial{\bar{\partial}}(\rho g_{2})=\int_{C}\rho g_{1}{\bar{\partial}}\partial(\rho g_{2})=\int_{C}d(\rho g_{1}\partial\rho g_{2})-\int_{C}{\bar{\partial}}(\rho g_{1})\wedge\partial(\rho g_{2})=$
@@
We have
@@
The last sentence follows since if $\rho g_{i}$ are both $C^{\infty}$, clearly $f_{i}g_{1}d(f_{i}g_{2})$ has no poles on $Z$.
@@ 4. Second fundamental form and Hodge Gaussian map
Denote by ${\mathcal{M}}_{g}$ the moduli space of smooth complex projective curves of genus $g\geq 4$ and let ${\mathcal{A}}_{g}$ be the moduli space of principally polarised abelian varieties of dimension $g$. The space ${\mathcal{A}}_{g}$ is a quotient of the Siegel space ${\mathcal{H}}_{g}=Sp(2g,{\mathbb{R}})/U(g)$ by the action of the symplectic group $Sp(2g,{\mathbb{Z}})$. The space ${\mathcal{H}}_{g}$ is a Hermitain symmetric domain and it is endowed with a canonical symmetric metric. We consider the induced orbifold metric (called the Siegel metric) on the quotient ${\mathcal{A}}_{g}$.
@@
the Torelli map, where $j(C)$ is the Jacobian of $C$ and $\Theta_{C}$ is the principal polarisation induced by cup product. It is an orbifold embedding outside the hyperelliptic locus ([25]).
@@
In [5, Theorem 2.1] it is proven that $II$ is equal (up to a constant) to the Hodge-Gaussian map $\rho$ of [5, Proposition-Definition 1.3]. We briefly recall its definition.
Let $Q=\sum_{i,j}a_{i,j}\omega_{i}\odot\omega_{j}\in I_{2}$, where $\omega_{i}\in H^{0}(K_{C})$, and $\Theta\in{\mathcal{A}}^{0,1}(T_{C})$, $[\Theta]=:\zeta.$
@@
(6) $II(\sum_{i,j}a_{ij}\omega_{i}\odot\omega_{j})(\zeta)=\sum_{i,j}a_{ij}\omega_{i}\partial h_{j}.$
In [5, Theorem 3.1] (see also [3, Theorem 2.2]) it is proven that if $C$ is a non-hyperelliptic curve of genus $g\geq 4$ and $p,q$ are two distinct points in $C$, we have
(7) $II(Q)(\xi_{p},\xi_{p})=-2\pi i\cdot\mu_{2}(Q)(p),$
@@
Since $II$ is injective (see [4, Corollary 3.4]), $II(I_{2})\subset Sym^{2}H^{1}(T_{C})^{\vee}$ is a linear system of quadrics in ${\mathbb{P}}H^{1}(T_{C})\cong{\mathbb{P}}^{3g-4}$ of dimension $\frac{(g-2)(g-3)}{2}$. Hence for every curve $C$ of genus $g\leq 9$, $dim(II(I_{2}))<3g-4$, so the intersections of the quadrics in $II(I_{2})$ is non empty, thus there exist asymptotic directions.
@@
Fix a holomorphic form $\omega\in H^{0}(K_{C}),$ $\omega\neq 0,$ with zero locus $Z$. With the notation of section 2, we set $\zeta=[{\bar{\partial}}(\frac{\rho}{\omega})]$. So we have $\zeta\omega=0$ and $\Theta\omega_{j}=\gamma_{j}+{\bar{\partial}}h_{j}$, $\gamma_{j}$ harmonic. Let $g_{i}$ be the meromorphic function given by $\frac{\omega_{i}}{\omega}$. We would like compute $II(Q)(\zeta\odot\zeta).$
@@
So, using Lemma (3.2) we get
$II(Q)(\zeta\odot\zeta)=-\sum_{i=1}^{s}\int_{C}(\partial h_{2i}\wedge{\bar{\partial}}h_{2i-1}+\partial h_{2i-1}\wedge{\bar{\partial}}h_{2i})=$
$-\sum_{i=1}^{s}[(w(\zeta,\omega_{2i-1},\omega_{2i})+\int_{C}\partial(\rho g_{2i})\wedge{\bar{\partial}}(\rho g_{2i-1}))+\int_{C}\partial(\rho g_{2i-1})\wedge{\bar{\partial}}(\rho g_{2i})].$

### v2

@@ 1. Introduction
Clearly a tangent direction to a totally geodesic subvariety is asymptotic. But the locus of asymptotic directions in the projective tangent bundle of the moduli space of curves ${\mathcal{M}}_{g}$ is a natural locus, that is worthwhile investigating. For example we believe that asymptotic directions could also be useful in the study of fibred surfaces in relation with Xiao conjecture ([35], [17], see Remark 4.3).
Since $II$ is injective (see [8, Corollary 3.4]), its image in $Sym^{2}H^{1}(T_{C})^{\vee}$ is a linear system of quadrics in ${\mathbb{P}}H^{1}(T_{C})\cong{\mathbb{P}}^{3g-4}$ of dimension $\frac{(g-2)(g-3)}{2}$. Hence for every curve $C$ of genus $g\leq 9$, $dim(II(I_{2}))<3g-4$, so the intersections of the quadrics in $II(I_{2})$ is non empty, thus there exist asymptotic directions. In fact for $g\leq 7$ there are examples of special subvarieties of ${\mathcal{A}}_{g}$ generically contained in the Torelli locus for $g\leq 7$ (see [11], [12], [13], [14], [27], [28], [29], [32], [33]).
@@
(Theorem 5.5, Theorem 5.1) Let $C$ be a smooth curve of genus $g\geq 4$, take an integer $d<Cliff(C)$ and an infinitesimal deformation $\zeta\in H^{1}(T_{C})$ of rank $d$. Then we have:
@@
as the locus of deformations of rank at most $d$. When $d=1$ it is the bicanonical curve.
In particular, if a curve $C$ of genus $g$ has maximal Clifford index, equal to $\lfloor\frac{g-1}{2}\rfloor$, this gives a characterisation of the image of the natural map ${\mathbb{P}}T_{C_{d}}\to{\mathbb{P}}(H^{1}(T_{C}))$ as the locus of deformations of rank at most $d$, for all $d<\lfloor\frac{g-1}{2}\rfloor$. We recall that this applies to the general curve in moduli space, which has maximal Clifford index. Moreover, for such values of $d$, Theorem 1.1(2), for curves of maximal Clifford index the base locus of the linear system of quadrics $II(I_{2})$ in ${\mathbb{P}}H^{1}(T_{C})$ does not contain any point $[\zeta]$ with $Rank(\zeta)\leq d$.
@@
More precisely, the linear combinations of Schiffer variations which are asymptotic in the above Theorem are $\xi_{p}\pm i\xi_{\sigma(p)}$, where $\sigma$ is the bielliptic involution and $(p,\sigma(p))\in C\times C$ is in the zero locus of the meromorphic form $\hat{\eta}\in H^{0}(K_{C\times C}(2\Delta))$ which determines the second fundamental form $II$ (see [7, Theorem 3.7]).
@@
In Section 4 we recall the definition of the second fundamental form, we define asymptotic directions and we use the results in Section 3 to give a formula that computes the second fundamental form $II(Q)(\zeta\odot\zeta)$ on some $\zeta$ with non trivial kernel, in terms of residues of meromorphic forms (see Proposition 4.4).
In Section 5 we consider infinitesimal deformations of rank $d$ less than the Clifford index of the curve, and we analyse the extension corresponding to $\zeta$ and the rank 2 vector bundle $E$ of the extension. First we show that if either $d<Cliff(C)$, or $d=Cliff(C)<\frac{g-1}{2}$ and $E$ is not globally generated, then $\zeta$ is a linear combination of Schiffer variations supported on a divisor $D$ of degree $d$ (see Theorem 5.1). One of the main technical tool is a Theorem of Segre-Nagata and Ghione (see [24] p. 84) on the existence of a subline bundle $A$ of $E$ such that $deg(A)\geq\frac{g-1}{2}$. Then we show that if $d<Cliff(C)$, no linear combinations of Schiffer variations of rank $d$ is asymptotic (see Theorem 5.4). Here we use a result of Green and Lazarsfeld ([19, Theorem 1]) that allows us to find some quadrics where our technique to compute the second fundamental form works well. Finally we prove Theorem 5.5.
@@ 2. Preliminaries on deformations
We want now to give an explicit description in Dolbeault cohomology of the image of $\delta$.
@@
Let $s\in H^{0}({\mathcal{O}}_{Z})$ be a holomorphic section and assume that $f_{i}(z_{i})=\sum_{j=0}^{n_{i}-1}\beta_{j,i}z_{i}^{j}$ is its polynomial expression in $U_{i}$. Let $\tilde{\rho}$ be a ${\mathcal{C}}^{\infty}$ function on $C$ which is equal to $1$ on $\cup_{i}D_{i}$ and equal to 00 on $C\setminus\cup_{i}\Delta_{i}.$
@@
Then, in Dolbeault cohomology we have $\delta(s)=[{\bar{\partial}}(\frac{\rho}{\omega})]$. Setting $\zeta:=[{\bar{\partial}}(\frac{\rho}{\omega})]$, clearly $\omega\cup\zeta=0$. Conversely, by the exact sequence (2), any $\zeta$ such that $\omega\cup\zeta=0$ is in the image of $\delta$, hence it has a Dolbeault representative as above.
@@
We say that a nonzero infinitesimal deformation $\zeta\in H^{1}(T_{C})$ is split if the vector bundle $E$ in (1) splits as a direct sum of line bundles $E=(K_{C}\otimes L^{\vee})\oplus L$.
@@
Let $C$ be a smooth complex projective curve of genus $g\geq 2$. Let $p$ be a point in $C$ and $z$ a local coordinate centred in $p$. We define the $n^{th}$ Schiffer variation at $p$ to be the element $\xi_{p}^{n}\in H^{1}(C,T_{C})\cong H^{0,1}_{\bar{\partial}}(T_{C})$ whose Dolbeault representative is ¯$\frac{\bar{\partial}\rho_{p}}{z^{n}}\frac{\partial}{\partial z}$, where $\rho_{p}$ is a bump function in $p$ which is equal to one in a small neighborhood $U$ containing $p$, $\xi_{p}^{n}=[\frac{\bar{\partial}\rho_{p}}{z^{n}}\frac{\partial}{\partial z}]$. Clearly $\xi_{p}^{n}$ depends on the choice of the local coordinate $z$. Take the exact sequence
@@
More generally, any linear combination $\zeta=\sum_{i=1}^{k}\sum_{j=1}^{m_{i}}b_{j,i}\xi^{j}_{p_{i}}$, with $\sum_{i=1}^{k}m_{i}\leq g$ has rank $\leq\sum_{i=1}^{k}m_{i}$, since $H^{0}(K_{C}(-\sum_{i=1}^{k}m_{i}p_{i}))\subset ker(\cup\zeta)$.
@@ 3. Some computations
Consider two holomorphic one forms $\omega_{1}$ and $\omega_{2}$. With the notation introduced in section 2, take an infinitesimal deformation $\zeta$ having a form $\omega\neq 0$ in its kernel and denote by $Z=\sum n_{i}p_{i}$ its zero divisor . Then $\zeta=[\Theta]\in H^{1}(T_{C})$ with $\Theta:={\bar{\partial}}(\frac{\rho}{\omega})$, and $\rho=\sum_{i}\tilde{\rho}f_{i}$. We have $\Theta\omega_{i}=\gamma_{i}+{\bar{\partial}}h_{i}$, where $\gamma_{i}$ is harmonic and $h_{i}$ is a ${\mathcal{C}}^{\infty}$ function on $C$, so $\zeta\cup\omega_{i}=[\gamma_{i}]\in H^{1}({\mathcal{O}}_{C})$. Consider the two meromorphic functions $g_{1}=\frac{\omega_{1}}{\omega}$ and $g_{2}=\frac{\omega_{2}}{\omega}$. Then clearly, by the above construction we have
@@
Note that if $\rho g_{1}$ is $C^{\infty}$, then $\omega_{1}\in ker(\cup\zeta)$ and $\gamma_{1}=0$. Moreover, up to a constant, we have $h_{1}=\rho g_{1}.$
@@
Since $\int_{C}\partial h_{1}\wedge\gamma_{2}=\int_{C}d(h_{1}\gamma_{2})=0$, and $\Theta\omega_{2}=\bar{\partial}(\rho g_{2})$, we get
@@
$=\int_{C}\partial(\rho g_{1})\wedge{\bar{\partial}}(\rho g_{2}),$
hence the first equality is proven. Moreover,
$\int_{C}\partial(\rho g_{1})\wedge{\bar{\partial}}(\rho g_{2})=\int_{C}d(\rho g_{1})\wedge{\bar{\partial}}(\rho g_{2})=\int_{C}d(\rho g_{1}{\bar{\partial}}(\rho g_{2}))-\int_{C}\rho g_{1}\partial({\bar{\partial}}(\rho g_{2})).$
Using Stokes theorem, and recalling that ¯${\bar{\partial}}(\rho g_{2})=\Theta\omega_{2}$ is a ${\mathcal{C}}^{\infty}$ $(0,1)$-form on $C$, we have
$\int_{C}d(\rho g_{1}{\bar{\partial}}(\rho g_{2}))=0.$
Then we get
$\int_{C}\partial h_{1}\wedge{\bar{\partial}}h_{2}=-\int_{C}\rho g_{1}\partial{\bar{\partial}}(\rho g_{2})=\int_{C}\rho g_{1}{\bar{\partial}}\partial(\rho g_{2})=\int_{C}d(\rho g_{1}\partial(\rho g_{2}))-\int_{C}{\bar{\partial}}(\rho g_{1})\wedge\partial(\rho g_{2})=$
@@
Now, taking $D_{i,\epsilon}$ a small disc around $p_{i}$ where $\rho=f_{i}$, and by Stokes theorem, we have
@@
So we have shown that
$\int_{C}\partial h_{1}\wedge{\bar{\partial}}h_{2}=\int_{C}\partial(\rho g_{2})\wedge{\bar{\partial}}(\rho g_{1})+w(\zeta,\omega_{1},\omega_{2}),$
and by Remark 3.1 the first term is equal to $\int_{C}\partial h_{2}\wedge\overline{\partial}h_{1}.$
Finally, if $\rho g_{1}$ and $\rho g_{2}$ are both $C^{\infty}$, clearly $f_{i}g_{1}d(f_{i}g_{2})$ has no poles on $Z$, hence the last sentence follows.
@@ 4. Second fundamental form and Hodge Gaussian map
Denote by ${\mathcal{M}}_{g}$ the moduli space of smooth complex projective curves of genus $g$ and let ${\mathcal{A}}_{g}$ be the moduli space of principally polarised abelian varieties of dimension $g$. The space ${\mathcal{A}}_{g}$ is a quotient of the Siegel space ${\mathcal{H}}_{g}=Sp(2g,{\mathbb{R}})/U(g)$ by the action of the symplectic group $Sp(2g,{\mathbb{Z}})$. The space ${\mathcal{H}}_{g}$ is a Hermitain symmetric domain and it is endowed with a canonical symmetric metric. We consider the induced orbifold metric (called the Siegel metric) on the quotient ${\mathcal{A}}_{g}$.
@@
the Torelli map, where $j(C)$ is the Jacobian of $C$ and $\Theta_{C}$ is the principal polarisation induced by cup product. If $g\geq 3$, it is an orbifold embedding outside the hyperelliptic locus ([30]).
@@
From now on we will assume $g\geq 4$, so $N_{{\mathcal{M}}_{g}^{0}/{\mathcal{A}}_{g}}^{*}$ is non trivial.
@@
In [9, Theorem 2.1] it is proven that $II$ is equal (up to a constant) to the Hodge-Gaussian map $\rho$ of [9, Proposition-Definition 1.3]. We briefly recall its definition.
Let $Q=\sum_{i,j}a_{i,j}\omega_{i}\otimes\omega_{j}\in I_{2}$, with $a_{i,j}=a_{j,i}$, $\omega_{i}\in H^{0}(K_{C})$, and $\Theta\in{\mathcal{A}}^{0,1}(T_{C})$, $[\Theta]=:\zeta.$
@@
(6) $II(\sum_{i,j}a_{ij}\omega_{i}\otimes\omega_{j})(\zeta)=\sum_{i,j}a_{ij}\omega_{i}\partial h_{j}.$
In [9, Theorem 3.1] (see also [7, Theorem 2.2]) it is proven that if $C$ is a non-hyperelliptic curve of genus $g\geq 4$ and $p,q$ are two distinct points in $C$, we have
(7) $II(Q)(\xi_{p}\odot\xi_{p})=-2\pi i\cdot\mu_{2}(Q)(p),$
@@
Since $II$ is injective (see [8, Corollary 3.4]), $II(I_{2})\subset Sym^{2}H^{1}(T_{C})^{\vee}$ is a linear system of quadrics in ${\mathbb{P}}H^{1}(T_{C})\cong{\mathbb{P}}^{3g-4}$ of dimension $\frac{(g-2)(g-3)}{2}$. Hence for every curve $C$ of genus $g\leq 9$, $dim(II(I_{2}))<3g-4$, so the intersections of the quadrics in $II(I_{2})$ is non empty, thus there exist asymptotic directions.
@@
Remark 4.3.
We recall that the original Xiao conjecture for non isotrivial fibered surfaces $S\to B$ with general fibre $F$ of genus $g$ says that the relative irregularity $q_{f}$ safisfies the inequality $q_{f}\leq\frac{g+1}{2}$. In [31], [2] counterexamples to this conjecture have been given, and a modified version, $q_{f}\leq\left\lceil{\frac{g+1}{2}}\right\rceil$ is stated in [4]. In two of the four counterexamples, the associated modular map $p:B\to M_{g}$ gives rise to curves contained in a totally geodesic subvariety ([14], [23], [33]). Moreover this always happens when $q_{f}=g-1.$ In fact in that case there is a fixed part $A\times B\to B$ of the Jacobian fibration $J\to B$ where $A$ is an abelian variety of dimension $g-1.$ Then the period variation of the Jacobian fibration is only given by the elliptic fibration given by the quotient $J/A.$ So the image of the period map $B\to A_{g}$ is a totally geodesic curve. Since the tangent vectors to a totally geodesic curve are asymptotic directions, these fibrations give rise to asymptotic directions whose rank is $\leq g-q_{f}.$ So the vanishing of the second fundamental form on tangent directions give new infinitesimal constraints that could help to classify fibered surfaces with $q_{f}>\frac{g+1}{2}$.
@@
Fix a holomorphic form $\omega\in H^{0}(K_{C}),$ $\omega\neq 0,$ with zero locus $Z$. With the notation of section 2, we set $\zeta=[{\bar{\partial}}(\frac{\rho}{\omega})]$. So we have $\zeta\omega=0$ and $\Theta\omega_{j}=\gamma_{j}+{\bar{\partial}}h_{j}$, $\gamma_{j}$ harmonic. Let $g_{i}$ be the meromorphic function given by $\frac{\omega_{i}}{\omega}$. We would like to compute $II(Q)(\zeta\odot\zeta).$
@@
By Remark 3.1 we have
$II(Q)(\zeta\odot\zeta)=-\sum_{i=1}^{s}\int_{C}(\partial h_{2i}\wedge{\bar{\partial}}h_{2i-1}+\partial h_{2i-1}\wedge{\bar{\partial}}h_{2i})=-2\sum_{i=1}^{s}\int_{C}(\partial h_{2i-1}\wedge{\bar{\partial}}h_{2i}).$
Using Lemma 3.2 we have
$\int_{C}(\partial h_{2i-1}\wedge{\bar{\partial}}h_{2i})=\int_{C}\partial(\rho g_{2i-1})\wedge{\bar{\partial}}(\rho g_{2i})=w(\zeta,\omega_{2i-1},\omega_{2i})+\int_{C}\partial(\rho g_{2i})\wedge{\bar{\partial}}(\rho g_{2i-1}).$
So we obtain
$II(Q)(\zeta\odot\zeta)=-\sum_{i=1}^{s}[(w(\zeta,\omega_{2i-1},\omega_{2i})+\int_{C}\partial(\rho g_{2i})\wedge{\bar{\partial}}(\rho g_{2i-1}))+\int_{C}\partial(\rho g_{2i-1})\wedge{\bar{\partial}}(\rho g_{2i})].$